solve for x.
3x + 1 +6x - 5 + 110=180

Answer by nyc_function(2733)

3x + 1 + 6x - 5 + 110 = 180
9x + 106 = 180
9x = 180 - 106
9x = 74
x = 74/9
x = 8.2